OpenStudy (anonymous):

find the fourth term of the geometric sequence whose first term is 6 and whose common ratio is -5

OpenStudy (anonymous):

use the explicit definition\[a(n)=a _{1}r^{n-1}\]where a1=6; r=-5, and n=4\[a(4)=6*(-5)^{4-1}=-750\]